40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13328     Entry time: Fri Sep 22 18:12:27 2017     In reply to: 13315     Reply to this: 13336
Author: gautam 
Type: Update 
Category: LSC 
Subject: DAC noise measurement (again) 

I've been working on setting up some scripts for measuring the DAC noise.

In all the DRMI noise budgets I've posted, the coil-driver noise contribution has been based on this measurement, which could be improved in a couple of ways:

  • The measurement was made at the output of the AI board - we can make the measurement at the output of the coil driver board, which will be a closer reflection of the actual current noise at the OSEM coils.
  • The measurement was made by driving the DAC with shaped random noise - but we can record the signal to the coils during a lock and make the noise measurement by using awg to drive the coil with this signal, with elliptic bandstops at appropriate frequencies to reveal the electronics noise.
  1. The IN1 signals to the coils aren't DQ-ed, but ideally this is the signal we want to inject into the coil_EXC channel for this measurement - so I re-locked the DRMI a couple of nights ago and downloaded the coil IN1 channel data for ~5mins for the ITMs and BS.
  2. AWGGUI supposedly has a feature that allows you to drive an EXC channel with an arbitrary signal - but I couldn't figure out how to get this working. I did find some examples of this kind of application using the Python awg packages, so I cobbled together some scripts that allows me to drive some channels and place elliptic bandstop filters as I required. 
  3. I wasted quite a bit of time trying to implement these signals in Python using available scipy functions, on account of me being a DSP n00b frown. When trying to design discrete-time filters, of course numerical precision errors become important. Initially I was trying to do everything in the "Transfer function (numerator-denominator)" basis, but as Rana pointed out, the way to go is using SOSs. Fortunately, this is a simple additional argument to the relevant python functions, after which elliptic bandstop filter design was trivial.
  4. The actual test was done as follows:
    • Save EPICS PIT/YAW offsets, set them to 0, disable Oplev servos, and then shut down optic watchdog once the optic is somewhat damped. This is to avoid the optics getting a large kick when disconnecting the DB15 connector from the coil driver board output.
    • Disconnect above-mentioned DB15 connector from the appropriate coil driver board output.
    • Turn off inputs to coils in filter module EPICs screens. Since the full signal (local damping servo output + Oplev servo output + LSC servo output) to the coil during a DRMI lock will be injected as an excitation, we don't need any other input. 
    • Use scripts (which I will upload to a git repo soon) to set up the appropriate excitation.
    • To measure the spectrum, I used a DB15 breakout board with test-points soldered on and some mini-grabber-to-BNC adaptors, in order to interface the SR785 to the coil driver output. We can use the two input channels of the SR785 to simultaneously measure two coil driver board output channels to save some time.
    • Take a measurement of the SR785 noise (at the appropriate "Input Range" setting) with inputs terminated to estimate the analyzer noise floor.
    • Just for kicks, I made the measurement with the de-whitening both OFF/ON.

I only managed to get in measurements for the BS and ITMX today. ITMY to be measured later, and data/analysis to follow.

The ITMX and BS alignments have been restored after this work in case anyone else wants to work with the IFO.


Some slow machine reboots were required today - c1susaux was down, and later, the MC autolocker got stuck because of c1iool0 being unresponsive. I thought we had removed all dependency of the autolocker on c1iool0 when we moved the "IFO-STATE" EPICS variable to the c1ioo model, but clearly there is still some dependancy. To be investigated.

ELOG V3.1.3-